Tuesday, 25 August

Tuesday, 25 August

Invited Lectures

\underline{14:00 – 14:45}

03 Loïc Merel, Université Paris 7 - Denis Diderot, Paris, France Determination of rational points using Dirichlet series (I: p.52) Chair: Wilhelm Zink

MA 005

08 Nicole Tomczak-Jaegermann, University of Alberta, Edmonton, Canada From finite- to infinite-dimensional phenomena in Geometric Functional Analysis, on local and asymptotic levels (I: p.108) Chair: W. Timothy Gowers

MA 001

12 Alain-Sol Sznitman, ETH Zürich, Zürich, Switzerland Brownian motion and random obstacles (I: p.149) Chair: Terry J. Lyons

H 1028

15 Percy Alec Deift, Courant Institute, New York University, New York, USA Uniform asymptotics for orthogonal polynomials (I: p.179) Chair: Tom H. Koornwinder

H 1012

16 Andreas W.M. Dress, Universität Bielefeld, Bielefeld, Germany The Tree of Life and Other Affine Buildings (I: p.186) Chair: Hanfried Lenz

H 2032

18 Mariolina Bartolini Bussi, Università di Modena, Modena, Italy Drawing instruments: the theory of a practice from history to didactics (I: p.197) Chair: Klaus-D. Graf

H 104

\underline{15:00 – 15:45}

!!!

Lecture by a Fields Medalist

— to be announced — Chair: N.N.

MA 001

03 Shinichi Mochizuki, RIMS, Kyoto University, Kyoto, Japan The Intrinsic Hodge Theory of p-adic Hyperbolic Curves (I: p.53) Chair: Helmut Koch

MA 005

12 Peter Gavin Hall, Australian National University, Canberra, Australia Applications of Intentionally Biased Bootstrap Methods (I: p.144) Chair: Olaf Bunke

H 1028

15 Jan-Olov Strömberg, Royal Institute of Technology, Stockholm, Sweden Apriori estimates for approximation by wavelets in higher dimensions (I: p.180) Chair: Peter Maaß

H 1012

16 Bonnie Berger, Massachusetts Institute of Technology, Cambridge, USA Sequence-based Methods for Protein Folding (I: p.185) Chair: Michiel Hazewinkel

H 2032

18 James Stigler, University of California, Los Angeles, USA Classroom Mathematics Instruction in Three Cultures: An Introduction to the TIMSS Video Study (I: p.200) Chair: Thomas Jahnke H 104 \underline{16:15 – 17:00}

01 Matthew Foreman, University of California, Irvine, USA Generic Large Cardinals: New Axioms for Mathematics? (I: p.30) Chair: Peter G. Koepke

H 1012

02 Aner Shalev, The Hebrew University, Jerusalem, Israel Simple groups, permutation groups, and probability (I: p.45) Chair: Bernhard H. Neumann

MA 005

05 Boris Dubrovin, International School for Advanced Studies, Trieste, Italy and Steklov Math. Institute, Moscow, Russia Geometry and analytic theory of Frobenius manifolds (I: p.69) Chair: Alexander I. Bobenko

MA 001

11 Eugene Bogomolny, Université Paris Sud, Orsay, France Spectral Statistics (I: p.129) Chair: Pierre Collet

H 1028

13 Neil J.A. Sloane, AT&T Research Labs, Florham Park, USA Recent developments on the sphere-packing problem (I: p.161) Chair: Martin Aigner

H 2032

\underline{16:15 – 18:00}

18 George Eyre Andrews, Pennsylvania State University, University Park, USA D.J. Lewis (moderator), NSF, Arlington, USA David Alexander Smith, Duke University, Durham, USA Collegiate Mathematics Instruction for the 21st Century — A Panel Discussion (I: p.200) Chair:

H 104

\underline{17:15 – 18:00}

01 Ludomir Newelski, University of Wrocław, Wrocław and Polish Academy of Sciences, Poland Meager forking and m-independence (I: p.33) Chair: Alex James Wilkie

H 1012

02 Jeremy C. Rickard, University of Bristol, Bristol, England The abelian defect group conjecture (I: p.45) Chair: Dieter Happel

MA 005

11 Horng-Tzer Yau, Courant Institute, New York University, New York, USA Scaling Limit of Particle Systems, Incompressible Navier-Stokes Equation, and Boltzmann Equation (I: p.138) Chair: Ruedi Seiler

H 1028

13 Joseph Adolphe François Thas, University of Gent, Gent, Belgium Finite Geometries, Varieties and Codes (I: p.161) Chair: Theo Grundhöfer

H 2032

16 Rolf Rannacher, Universität Heidelberg, Heidelberg, Germany Mathematics in Numerical Flow Simulation (I: p.192) Chair: Rolf Jeltsch MA 001